Let α = {1 + 2t, t − t 2 , t + t 2}

(a) Show that α is a basis for P2(R).

(b) Let p(t) = 1 + 3t + t 2 . Find [p(t)]α.

(c) Define the transformation T : P2(R) → P2(R) as T (p(t)) = p 0 (t) − p(t) i.e., the difference of p(t) and its first derivative. Determine whether this transformation is a linear transformation.

(d) Find [T]αProblem 4. Let a = {1 + 2t, t – t?,t+t²} [ 4 pts] (a) Show that a is a basis for P2 (R). (b) Let p(t) = 1 + 3t + t2. Find [p(
Answer #1

4) 9). d = { H2t, t-42 +62] Now show d is basis for PalR) (1+2+) +2 (t-t?) + 3 (H1750 (1+261 +62 +6372 +612 +(3) ť? 50 ci bru0 +05) (14351 - 2 trzy 283=2 so (23 I- C2 (351 -O so (pallax H Now Check o T! PER) + PSR) TfPct ) )= pkt)- ple) T Ů linear ba(2+35 -2-201 (2435-2-2 (24(3= -4 +0=) C2+(3= -4 2+(3=0 203=-4 from o C3=-2 -2 T((1+2+))s 120 (18+ k) (23)+(27-7) (ZH (+2+1) 2